# Two-center harmonic oscillator basis for Skyrme-DFT calculations (I): formalism and Proof of Principle

### v3

@@ I Introduction
In the last decades, we have seen two different approaches employing the two-center method. The first one relies on extending an external potential (usually Wood-Saxon) into a two-center version Scharnweber et al. (1971); Chandra and Mosel (1978) and solving it on a two-center basis; the harmonic oscillator (HO) being the usual choice. After this, the Schrödinger equation is solved using the Green’s function method Diaz-Torres (2008). Even though the results shown for both light Hagino and Ichikawa (2017); von Oertzen (1996) and heavy systems Mirea (2008) look promising, the Coulomb interaction is not accurately treated. Indeed, in some cases, it is just approximated as a charged spherical distribution, which does not treat overlapping fragments properly. The computational burden associated with obtaining the Green’s function for large basis sets, and the bad solvability conditions when s.p. crossings occur, make this approach optimal for light masses, but cumbersome when describing the evolution of heavy nuclei along the fission path.
@@
If we look for an approach more suitable for heavy nuclei such as DFT, the number of references decreases drastically. Since the pioneering work of Berger and Gogny Berger and Gogny (1980) there have been only a few applications of this technology to covariant DFT (CDFT) Geng et al. (2007); Li et al. (2024). However, these aforementioned implementations were restricted to axially deformed bases, which can be a handicap when trying to model the fission paths that are known to pass through triaxial shapes. It appears that only by controlling the deformation and orientations of the molecular-like components, along with the separation between them, we may see the real impact of the collectivity of the initial states in the post-scission configurations.
In the present work, we focus on the underpinnings of the two-center (Cartesian) HO basis method and its implementation suitable for the Skyrme energy density functional. We embedded the method in the newest version of the DFT solver hfodd Dobaczewski et al. (2021); J. Dobaczewski, et al. (2024), where the user now can set the deformation and separation of the bases as input data to perform the Hartree-Fock (HF) or HF-Bogoliubov calculations.
The structure of the paper is as follows. In section II we present the theoretical framework, highlighting how the usual one-center HO (OCHO) self-consistent procedure can be reformulated to be used within the two-center HO (TCHO) basis. In particular, we show how to treat the Coulomb interaction exactly, both in the direct and exchange channels. Appendices A–C present all details of the implementation. In sections III and IV, we discuss the results of simple calculations performed for small systems, namely, for 8Be, where the Coulomb interaction is tested, and for the symmetric fission channel of 24Mg. Both Proof-of-Princple TCHO calculations are compared against the standard OCHO results. In section V, we provide final remarks and discuss potential applications of this newly developed method in future research.
@@ II.3 Matrix elements of the Skyrme mean field
where are the 33 orthogonal rotation matrices for centers A an B. Therefore, for non-co-axial bases, the Gauss-Hermite integration lattices have to be built by diagonalising the quadratic form in the coordinates, which defines the Gaussian factor in Eq. (17), and then determining the polynomial factors on those lattices. The use of non-co-axial bases will be discussed in a forthcoming publication.

### v5

@@ I Introduction
In the last decades, we have seen two different approaches employing the two-center method. The first one relies on extending an external potential (usually Wood-Saxon) into a two-center version Scharnweber et al. (1971); Chandra and Mosel (1978) and solving it on a two-center basis; the harmonic oscillator (HO) being the usual choice. After this, the Schrödinger equation is solved using the Green’s function method Diaz-Torres (2008). Even though the results shown for both light Hagino and Ichikawa (2017); von Oertzen (1996) and heavy systems J. and W. (1972); Mirea (2008) look promising, the Coulomb interaction is not accurately treated. Indeed, in some cases, it is just approximated as a charged spherical distribution, which does not treat overlapping fragments properly. The computational burden associated with obtaining the Green’s function for large basis sets, and the bad solvability conditions when s.p. crossings occur, make this approach optimal for light masses, but cumbersome when describing the evolution of heavy nuclei along the fission path.
@@
In the pursuit of suitable approaches for heavy nuclei, Density Functional Theory (DFT) stands out, although only a few groups utilized the two-center HO (TCHO) basis expansion. Since the pioneering work of Berger and Gogny more than 40 years ago Berger and Gogny (1980), a handful of axial applications for the Gogny functional exist in areas such as mass distributions Dubray et al. (2008); Regnier et al. (2016, 2019), fission barriers, and spontaneous fission lifetimes Berger et al. (1989); Lemaître et al. (2018). However, the Gogny implementation has never been published, leaving its applicability largely unknown. In addition, the Coulomb interaction was approximated by that of point-like charges near the scission point Berger et al. (1989), which may be inadequate for accurately describing the angular momentum generation of fission fragments. In the realm of Covariant DFT (CDFT), more details are available on the TCHO basis method Geng et al. (2007); Li et al. (2024), although this approach also relies on axially deformed co-axial basis states. So far no implementation of the general framework of the triaxial, shifted, and non-co-axial TCHO basis expansions exists. While this strategy has been explored in molecular physics Maintz et al. (2016), and some analytical expressions can be applied to the two-center expansion Nazmitdinov et al. (1996), there is no available nuclear DFT code that exploits this formalism.
In this work, we focus on the fundamentals of the two-center (Cartesian) HO basis method and its implementation suitable for the Skyrme energy density functional. We provide detailed information on how to compute the TCHO matrix elements relying on numerical integration instead of transformation of coefficients Robledo (2022) and how to address the generalized eigenvalue problem in the Hartree-Fock (HF) and HF-Bogoliubov (HFB) cases. This method has been integrated into the latest version of the DFT solver hfodd Dobaczewski et al. (2021); J. Dobaczewski, et al. (2024), where users can specify the deformation and separation of the bases as input data to perform the HF or HFB calculations.
The structure of the paper is as follows. In section II we present the theoretical framework, highlighting how the usual Skyrme+Coulomb one-center HO (OCHO) self-consistent procedure can be reformulated to be used within the TCHO basis. In particular, we show how to treat the Coulomb interaction exactly, both in the direct and exchange channels. Appendices A–C present all details of the implementation. In sections III and IV, we discuss the results of simple calculations performed for small systems, namely, for 8Be, where the Coulomb interaction is tested, and for the symmetric fission channel of 24Mg. Both Proof-of-Princple TCHO calculations are compared against the standard OCHO results. In section V, we provide final remarks and discuss potential applications of this newly developed method in future research.
@@ II.3 Matrix elements of the Skyrme mean field
where are the 33 orthogonal rotation matrices for centers A an B.
However, even for non-co-axial bases, the integration lattices remain the same, as the exponent of the common Gaussian is invariant under rotations. Thus, the expressions shown within this section and the appendices are still valid. The use of non-co-axial bases will be discussed in a forthcoming publication.